# Explainable Malware Analysis: Concepts, Approaches and Challenges

### v1

@@ IV-B Android-Based Malware Approaches
TABLE III: Research Addressing Explainable Machine Learning in Android Malware
@@
Kumar et al. 2018  [60] Convert high-dimensional data into low-dimensional data, which reduces the dimension and avoids extracting unnecessary features Introduced Static analysis and Feature vectorization methodologies for the analysis of Android malware Limited to smaller labeled datasets Refined to only static approach, should use hybrid and online approaches to examine the dynamic features and source code in run-time Feature Extraction
Melis et al. 2018  [80] Emphasizes the global characteristics learned by the model to distinguish between malicious and benign apps Provided general approach to explain the black-box Android malware detection Used gradient-based approach to identify whether a sample is correctly identified as malware or not by its most influential local features Provided interpretable decisions Approach to nonlinear model for Android malware No explanation for how various surrogate models affect the provided explanations. Global Surrogate
Iadarola et al. 2021  [49] Effectiveness of malware class extraction through deep learning approach Proposed a deep learning model for malware family identification Used Grad-CAM technique to predict the classes by detecting the important regions in the image Approach is vulnerable to Obfuscation techniques Grad-CAM
Kinkead et al. 2021  [56] Compare and analyze the similarities of the opcode sequence location, which is assumed malicious by CNN, to that of LIME Proposed a CNN-based novel method to identify the opcode sequence locations that are assumed to be malicious Their model achieved high accuracy, which states that CNN accomplishes exceptional performance with the DREBIN data set Lacking explanation in Filter activations LIME
Lu and Thing 2021  [71] To detect Android malware, the authors focused on the features attribution problem Optimization of features attribution by minimizing the variance of prediction scores and assigning high attribution values They made use of LIME and SHAP to compare the explanation ability of MPT explainer Mitigated methods for attack are not presented Limited to only one specific attack Modern Portfolio Theory (MPT)
Korine and Hendler 2021  [57] Explainable Model agnostic malware classification, which is both dataset and platform agnostic Presented effective and explainable malware classifier DAEMON Used Layer-wise Propagation explainable technique to explain the model DAEMON achieved high classification accuracy, and classification results are explainable Lack of evaluation of the ability of DAEMON classifier in order to classify the benign executables Feature Importance
Scalas 2021  [101] Identifying and characterizing effective Android ransomware detection Identifying traits that result in characterizing effective ransomware detection To assess the validity of features, explainability techniques are used to propose methods This work explored the relationship between explainable ML and adversarial attacks Missing explanation regarding usage of attribution technique Integrated Gradient
Yan et al. 2021  [125] Developing a novel method to extract rules from DNN rule extraction and facilitates its explainability Proposed a novel method for extracting rules from DNN, which ensures high accuracy Used extracted rules to identify malicious network traffic To detect mobile malware in an increased network environment, this work proposed an online detection system No proper explanation on the mentioned online detection system Rule-Based Learner
Wang et al. 2016  [118] This work focused on effectively using network traffic in the detection of mobile malware and explained the detection result using explainable machine learning techniques Proposed TrafficAV to detect malware traffic Extracted the features and combined them with a machine learning detection algorithm to detect the malicious application Limited to smaller malware samples Limited analysis of detection models Feature Extraction
Iadarola et al. 2021  [50] This work focused on The development of an explainable deep learning algorithm for mobile malware classification and detection Used the Grad-CAM explainability approach to describe mobile analysis in terms of images Localized the prominent portion of the image beneficial to the model in order to generate a specific accuracy Biased while labeling the data samples Grad-CAM
TABLE IV: Research Addressing Explainable Machine Learning in Android Malware (Continued).
Paper/Year Focus/Objective Contribution Limitation XAI technique
Wu et al. 2021  [122] This work focused on the classification of malware and interpreting the classification results of Android malware Proposed a novel interpretable approach (XMAL) by coupling it with an MPT explainer For interpreting with high fidelity, they generated the malware feature descriptions Conducted an online survey and qualitative analysis and provided a comparison of state-of-art approaches XMAL lacks the multi-attention mechanism Features are not sufficient enough to interpret the malicious behaviors XMAL, LIME
Alenezi and Ludwig 2021  [6] This work aims to use SHAP to explain the cybersecurity threats to data ML models such as Random Forest, XGBoost, and sequential model and applied SHAP explainability technique to explain it further They showed that XGBoost outperformed the other models by correctly classifying the highest no. of samples Missing proof that other ML models may/may not gain superior performance SHAP
Ullah et al. 2022  [115] This work aims to utilize SHAP to explain the important features for transfer learning and graphical malware features to detect malware Proposed android malware detection system with BERT-based transfer learning Utilized various ensembled machine-learning algorithms for malware classification and detection Utilized SHAP to explain the most important features that are responsible for the model’s outcome Missing results of Local Interpretable Model-agnostic Explanation (LIME) method SHAP explanations are limited to local scope SHAP
Naeem et al. 2022  [86] This work aims to provide explainable malware detection using image visualization and fine-tune the CNN-based transfer learning model Proposed fine-tuning CNN-based model for IoT device malware classification Utilized various machine-learning algorithms to test the effectiveness of IoT device malware classification Grad-CAM was utilized to develop heatmaps to assist security analysts in better understanding the classifier selection There is some reliability. However, thorough validation is lacking. Furthermore, it remains to be seen whether their framework can be used for large malware variants Grad-CAM
Alani and Awad. 2022  [4] This work aims to propose an explainable and accurate framework for identifying Android malware Performed feature selection for selecting the crucial features that can provide high accuracy Explaining the importance of selected features using the SHAP explainability technique Performing feature reduction and creating a reduced dataset with 35 effective features that can be used for Android malware detection SHAP explainability is restricted to only global scope SHAP
Liu et al. 2022  [70] Investigating the performance of ML models under realistic and unrealistic experimental setups Identifying the temporal inconsistency problem in Android malware detection Leveraging XAI to realize the reason behind the high performance of ML-based detection techniques The XAI techniques focus on interpretability from a feature-importance perspective that may not fully capture the complexity of model decisions, especially in DL-based models where interactions between features can be non-linear and more opaque. XMAL, DREBIN, and Model-Agnostic Techniques
Ambekar et al. 2024  [8] The paper focuses on enhancing Android malware classification by integrating attention mechanisms with LSTMs and explainable AI techniques to improve accuracy and interpretability. The paper introduces a novel and interpretable framework called TabLSTMNet that utilizes recent datasets to classify Android malware permissions. It employs XAI methodologies to elucidate the contributions of the features and enhance the transparency and effectiveness of the TabLSTMNet framework. The use of RandomOverSampler to address class imbalance by replicating instances of the minority class increases the likelihood of overfitting. The integration of TabNet and LSTM models, along with the XAI techniques, increases the computational complexity. LIME, and SHAP
Soi et al. 2024  [109] It enhances the explainability of Android malware detection systems by utilizing static analysis of function calls within the APK files. The paper introduces a new approach by selecting and analyzing API calls from the Function Call Graph (FCG) of Android applications. The paper employs SHAP values to provide clear, model-agnostic explanations for the classifier’s decisions. The large feature set (API calls) used for classification and explainability might pose challenges in terms of processing and interpretation. SHAP
@@
A standout innovation in this field is the DREBIN malware detection system [10], illustrated in Figure 9. DREBIN leverages a lightweight approach to identify Android malware on smartphones through static analysis, extracting application features that are represented in a binary vector format. This setup enables linear classification to differentiate between features of benign and malicious applications. Furthermore, DREBIN is distinguished by its explainable approach to malware detection. It provides insights into the reasoning behind its decisions by highlighting key attributes of detected malware.
@@
To analyze the Android malware, the research by Kumar et al. [60] introduces two ML-supported methodologies: one focuses on static analysis and the other on feature extraction. They perform feature extraction on the DREBIN dataset through vectorization, followed by feature selection and dimensionality reduction, thus transforming high-dimensional data into a more manageable low-dimensional format while omitting extraneous features. Their analysis yields metrics such as the True Positive Rate (TPR) and False Positive Rate (FPR), with their methodology demonstrating high precision and recall. Various ML-based algorithms, including SVM, KNN, Naive Bayes, and C4.5, are applied to the newly processed data, which is revealed SVM’s superior performance. The combination of static analysis, feature vectorization, and supervised learning enables these ML algorithms to identify new malware families with high true positive and recall rates. The subsequent subsection explores explainable analysis techniques for Android malware.
Figure 9: Schematic Representation of DREBIN malware detector [80].
@@ IV-B2 Model-Agnostic Based Approach
TABLE V: Research Addressing Explainable Machine Learning in Malware Analysis.
@@
Pan et al. 2020  [88] Developed regression-based machine learning model that uses linear regression to identify the major contributors amongst input features Proposed framework for malware detection using explainable Machine learning, which involves training, regression, and interpretation Filter activations are not well explained Limited analysis of interpretable techniques Linear regression
Pan et al. 2022  [89] Aims to focus on hardware-assisted malware detection and localization through Explainable AI Proposed method that provides an interpretable explanation to address the challenge of transparency of the classification results Showed the explainable result through efficient use of hardware performance counters (HPC) and embedded trace buffers (ETC) to provide precise localization of the malicious behavior. Lack of exploration of how different surrogate models can influence the results of their interpretation outcomes. Linear regression and Decision tree
Li et al. 2021  [64] Aims to provide an interpretable hardware malware detector to investigate the impact of feature contribution for a particular outcome Proposed I-MAD that understands the assembly code at the basic block at the executable level To provide the interpretations, they proposed Interpreted feed-forward neural network (IFFNN) I-MAD provides great prediction accuracy and understandable explainability and helps analysts in locating payloads and recurring patterns in malware samples There is no systemic description of the interpretation factors With their interpretable I-MAD model, adversaries have complete access to their model and can utilize the interpretations to generate adversarial samples IFFNN
@@
Wang et al. 2021  [117] Focuses on malware detection, which locates code snippets, and on effectively explaining the decision of malware classifier Proposed effective deep learning approach for Linux malware detection Methodology detects malicious behavior in malware that uses inline assembly Explanation of malware classification results using LRP Lacks clear future technological guidance Layer-wise relevance propagation (LRP)
Wang et al. 2021  [119] Exposing the weakness of malware detectors with explainability-guided evasion attacks Designed adversarial evasion attacks that depend on feature and problem-space manipulation Explanation of attacks they used model agnostic method - SHAP Missing proof that other explainable models may/may not gain superior performance Restricted to static malware detection Could not detect malicious behavior during run time SHAP
Mills et al. 2019  [83] Aims to provide interpretable real-time detection of malware using Random Forest (RF) Proposed an automated interpretable malware detection system called NODENS This system was able to detect the zero-day malware Limited to a smaller dataset Decision tree
@@
Kuppa and Le-Khac 2020  [61] Focused on taxonomy of black-box attacks for explainable AI methods that include a wide range of threat models and security properties in cyber security Proposed black-box approach that helps to design secure and explainable machine-learning methods Analyzed the accuracy, confidence, and consistency properties of gradient-based explainability methods Defensive mechanisms are not explained Gradient-Based Methods
Severi et al. 2021  [102] This work focused on backdoor poisoning attacks which influence the machine learning training process Proposed the use of explainable ML technique SHAP that creates effective backdoor triggers that are model agnostic, which influences the machine learning training process To preserve the binary’s functionality, the authors created a watermarking utility for Windows PE files Limited to specific attack methodologies (need to consider more generic attack) SHAP
Guo et al. 2018  [45] Aims to classify PDF malware and derive high-fidelity explanations for malware detection Designed a novel interpretability method LEMNA, which is a mixture of regression models and fused lasso In order to interpret the models, LEMNA considers the relevance and dependencies between the features According to the results, LEMNA performs better than LIME Explanations generated from the features are not addressed They paid less focus on comprehensive interpretability LEMNA
@@ IV-B3 Rule-Based Approach
Yan et al. [125] present an innovative approach for extracting rules from DNNs, aiming to balance the accuracy intrinsic to DNNs with the need for explainability in their operation. This methodology is visually summarized in Figure 10, which illustrates the comprehensive process of rule extraction and detection within a DNN framework. The initial phase contains the collection of network traffic data, utilizing a tool named DroidCollector for this purpose. Subsequent to data collection, feature extraction is conducted to distill the essential information necessary for training the model. This algorithm begins by verifying the appropriateness of the neural network settings, such as its suitability for classification tasks. In instances where the predicted label aligns with the true label, the model’s performance is deemed satisfactory. However, misclassification triggers a reassessment and update of the neural network’s weights. The backpropagation process prioritizes the weights of the outermost layer before sequentially addressing each subsequent layer, effectively distributing the errors from each output variable across the network’s hidden layers. Through numerous iterations, the model iteratively refines itself until it achieves optimal performance, at which point rules are extracted from the DNN.
Figure 10: DNN Rule Extraction and Detection [125].
@@
Also, as ML has emerged as a powerful tool for uncovering rules crucial for predictive data analysis, in their work, Wang et al. [118] develop TrafficAV an efficient and intelligible method for classifying mobile malware based on network traffic patterns. This approach is designed to minimize resource usage and performs malware detection and network traffic analysis server-side. TrafficAV leverages feature extraction combined with the C4.5 DT algorithm to detect the presence of malicious applications, applying two distinct detection models for HTTP and TCP protocols. This dual-model strategy has yielded high accuracy rates. Furthermore, TrafficAV provides an analysis of the significance of each feature in the decision-making process, offering user-friendly explanations of its findings.

### v2

@@ IV-A3 Image-Based Approach
By converting PE files into images, these techniques leverage CNNs’ strength in pattern recognition. Gradient-based visualization methods (e.g., Grad-CAM) then highlight which pixel regions most impact the model’s decision, increasing transparency for security analysts. The main challenge is the computational cost and the complexity of transforming binaries to images, but when accuracy and visual interpretability are desired, image-based XAI can be highly effective.
Collectively, these Windows-based XAI approaches illustrate a variety of explainability techniques—gradient-focused, model-agnostic, and image-based—each balancing interpretability, detection accuracy, and computational complexity. Next, we explore how Android malware detection demands similar yet distinct methods, given Android’s unique environment and features.
@@ IV-B Android-Based Malware Approaches
TABLE II: Research Addressing Explainable Machine Learning in Android Malware
@@
Kumar et al. 2018 [85] Reduce high-dimensional Android malware features to avoid unnecessary data Proposed static analysis + feature vectorization for effective Android malware detection Relies on small static datasets; lacks hybrid or dynamic analysis Feature Extraction
Melis et al. 2018 [86] Highlight global features to differentiate benign from malicious Android apps Presented a gradient-based, global surrogate method to explain black-box Android malware models Does not compare different surrogate models’ impact on explanations Global Surrogate
Iadarola et al. 2021 [87] Evaluate deep learning for Android malware family identification Introduced Grad-CAM-based CNN to localize crucial image regions for malware family classification Susceptible to code obfuscation Grad-CAM
Kinkead et al. 2021 [88] Compare CNN-identified malicious opcode sequences with LIME explanations Introduced a CNN to locate malicious opcode sequences; validated with LIME consistency No deep analysis of CNN filter activations LIME
Lu and Thing 2021 [89] Address Android malware detection via feature attribution Proposed MPT explainer to optimize feature attribution; compared with LIME/SHAP Limited scope on a single attack; no mitigation strategies Modern Portfolio Theory (MPT)
Korine and Hendler 2021 [90] Present DAEMON, a model-agnostic, explainable malware classifier Developed DAEMON with layer-wise propagation for explainable classification across platforms Unclear performance on benign executables Feature Importance
Scalas 2021 [91] Explore feature-based traits for effective Android ransomware detection Proposed a gradient-based approach and analyzed adversarial attacks with integrated gradients Lacks detailed rationale for specific attribution methods Integrated Gradient
Yan et al. 2021 [92] Introduce a rule-extraction method from DNNs for transparent mobile malware detection Extracts DNN rules for malicious traffic detection; offers high accuracy and explainability Lacks detail on the proposed online detection implementation Rule-Based Learner
Wang et al. 2016 [93] Leverage network traffic for mobile malware detection; add interpretable ML explanations Developed TrafficAV with feature extraction + ML classification for malware traffic Small dataset and limited model comparisons Feature Extraction
Iadarola et al. 2021 [94] Propose explainable DL for mobile malware detection and family classification Implemented Grad-CAM to highlight crucial image areas for accurate classification Potential label bias in the dataset Grad-CAM
Wu et al. 2021 [95] Combine MPT explainer (XMAL) with feature descriptions for high-fidelity Android malware classification Employs XMAL for interpretable detection, validated via user surveys and state-of-the-art comparison Missing multi-attention; limited feature scope XMAL, LIME
Alenezi and Ludwig 2021 [96] Apply SHAP for explainable cybersecurity threat detection Tested RF, XGBoost, and sequential models; SHAP reveals key features Limited model variety tested SHAP
Ullah et al. 2022 [97] Develop BERT-based transfer learning for Android malware, with SHAP for top features Merged text (BERT) and image-based (CNN) features; explained predictions with SHAP Only SHAP used, no LIME comparisons; local explanations only SHAP
Naeem et al. 2022 [98] Fine-tune CNN for IoT malware images, offering Grad-CAM-based explainability Applied Inception-v3 for IoT malware, employing Grad-CAM heatmaps for explanation Limited validation and scalability to larger malware families Grad-CAM
Alani and Awad. 2022 [99] Introduce PAIRED, a lightweight Android malware detector with SHAP feature importance Reduced features from 214 to 35, achieving  98% accuracy; SHAP explains global feature impact Only global SHAP used, no local interpretation SHAP
Liu et al. 2022 [100] Examine temporal bias in Android malware detection and use XAI to explain inflated accuracy Reveal how inconsistent time splits inflate ML performance, with XMal and Drebin for explanation Feature-importance XAI may miss deeper non-linear interactions XMAL, DREBIN, and Model-Agnostic Techniques
Ambekar et al. 2024 [101] Propose TabLSTMNet combining TabNet + LSTM for interpretable Android malware detection Achieves  97–98% accuracy, with LIME/SHAP explaining permission-level contributions Over-sampling risks overfitting; high computational complexity LIME, SHAP
Soi et al. 2024 [102] Use function-call graphs and SHAP for clear, model-agnostic explanations in Android Selects critical API calls from function-call graphs, employs SHAP for transparent classification Large-scale API features may hinder efficient analysis SHAP
@@
A standout innovation in this field is the DREBIN malware detection system [107]. DREBIN leverages a lightweight approach to identify Android malware on smartphones through static analysis, extracting application features that are represented in a binary vector format. This setup enables linear classification to differentiate between features of benign and malicious applications. Furthermore, DREBIN is distinguished by its explainable approach to malware detection. It provides insights into the reasoning behind its decisions by highlighting key attributes of detected malware.
@@
To analyze the Android malware, the research by Kumar et al. [85] introduces two ML-supported methodologies: one focuses on static analysis and the other on feature extraction. They perform feature extraction on the DREBIN dataset through vectorization, followed by feature selection and dimensionality reduction, thus transforming high-dimensional data into a more manageable low-dimensional format while omitting extraneous features. Their analysis yields metrics such as the True Positive Rate (TPR) and False Positive Rate (FPR), with their methodology demonstrating high precision and recall. Various ML-based algorithms, including SVM, KNN, Naive Bayes, and C4.5, are applied to the newly processed data, which reveals SVM’s superior performance. The combination of static analysis, feature vectorization, and supervised learning enables these ML algorithms to identify new malware families with high true positive and recall rates.
While feature extraction and dimensionality reduction can streamline malware detection, filtering out redundant instances in the dataset can further enhance both classifier efficiency and the clarity of explanations. Surendran et al. [108] propose using the Ochiai coefficient to identify and remove near-duplicate samples before retraining, which can reduce noise and training overhead. Eliminating these overlaps helps ensure that model explanations (e.g., SHAP or LIME outputs) capture truly informative patterns rather than repeated artifacts, ultimately improving both performance and the stability of XAI-based insights.
Following this discussion, below we will explore gradient-based approaches, which provide a more fine-grained analysis of feature importance in malware classification.
@@ IV-B1 Gradient-Based Approach
Gradient-based attribution helps Android malware detectors reveal exactly which app features (e.g., API calls, opcode sequences) most strongly affect the classifier’s output. While this granularity aids in pinpointing malicious behavior, these methods may require careful tuning to handle the massive variety of Android apps and can still be undermined by obfuscation or adversarial feature manipulation.
@@ IV-B2 Model-Agnostic Based Approach
Model-agnostic tools like LIME and SHAP allow security analysts to investigate any black-box Android malware detector, highlighting the top features—like permissions or API calls—responsible for a malicious label. Their interpretability fosters user trust but can sometimes yield inconsistent local explanations across app variants, especially if the underlying model is unstable.
TABLE III: Research Addressing Explainable Machine Learning in Malware Analysis.
@@
Pan et al. 2020 [110] Use linear regression for interpretable hardware-based malware detection Introduced a hardware trace–based detection framework with interpretable regression Insufficient details on filter/activation interpretability Linear regression
Pan et al. 2022 [111] Hardware performance counters + ETB for explainable malware detection/localization Provided interpretable HPC-based classification, localizing malicious behavior via RNN + DT Does not test various surrogates or compare interpretability outcomes Linear regression and Decision tree
Li et al. 2021 [112] Introduce I-MAD with an interpretable feed-forward neural network for hardware-level detection Analyzes assembly code, uses IFFNN to locate payload patterns and improve transparency Limited fidelity documentation; risk of adversarial misuse IFFNN
@@
Wang et al. 2021 [66] Pinpoint malicious code snippets in Linux malware with LRP-based explainability Proposed a CNN to detect inline assembly malware, using LRP to highlight influential code blocks Minimal discussion of future improvements Layer-wise relevance propagation (LRP)
Wang et al. 2021 [113] Develop SHAP-guided adversarial evasion on Linux malware detectors Combined feature-level and problem-space obfuscation using SHAP-based insights Limited to static detection and SHAP approach SHAP
Mills et al. 2019 [114] Propose NODENS, an interpretable RF-based real-time Linux malware detector Demonstrates zero-day detection with a tree-based approach for straightforward explanations Evaluation constrained by small dataset Decision tree
@@
Kuppa and Le-Khac 2020 [115] Explore black-box attacks on gradient-based XAI, using PDF and Android malware Investigated gradient-based XAI robustness and accuracy under black-box adversarial scenarios No detailed defense strategies were provided Gradient-Based Methods
Severi et al. 2021 [116] Demonstrate SHAP-based backdoor triggers that poison ML training Created stealthy backdoor triggers for PDF/Windows PE; validated with SHAP analysis Constrained to certain backdoor scenarios SHAP
Guo et al. 2018 [117] Introduce LEMNA for interpretable PDF malware classification with fused lasso LEMNA captures feature dependencies, outperforming LIME on PDF malware tasks Limited detail on feature-level explanation LEMNA
@@ IV-B3 Rule-Based Approach
Yan et al. [92] present an innovative approach for extracting rules from DNNs, aiming to balance the accuracy intrinsic to DNNs with the need for explainability in their operation. The initial phase contains the collection of network traffic data, utilizing a tool named DroidCollector for this purpose. Subsequent to data collection, feature extraction is conducted to distill the essential information necessary for training the model. This algorithm begins by verifying the appropriateness of the neural network settings, such as its suitability for classification tasks. In instances where the predicted label aligns with the true label, the model’s performance is deemed satisfactory. However, misclassification triggers a reassessment and update of the neural network’s weights. The backpropagation process prioritizes the weights of the outermost layer before sequentially addressing each subsequent layer, effectively distributing the errors from each output variable across the network’s hidden layers. Through numerous iterations, the model iteratively refines itself until it achieves optimal performance, at which point rules are extracted from the DNN.
@@
Also, as ML has emerged as a powerful tool for uncovering rules crucial for predictive data analysis, in their work, Wang et al. [93] develop TrafficAV, an efficient and intelligible method for classifying mobile malware based on network traffic patterns. This approach is designed to minimize resource usage and performs malware detection and network traffic analysis server-side. TrafficAV leverages feature extraction combined with the C4.5 DT algorithm to detect the presence of malicious applications, applying two distinct detection models for HTTP and TCP protocols. This dual-model strategy has yielded high accuracy rates. Furthermore, TrafficAV provides an analysis of the significance of each feature in the decision-making process, offering user-friendly explanations of its findings.
@@
Rule-based learners (e.g., DNN rule extraction or tree-based methods) strive to combine the accuracy of deep nets with the readability of logical rules. When applied to Android malware, these frameworks can yield straightforward if-then statements describing malicious behaviors. Still, as Android evolves, rule sets must be updated frequently, or attackers may exploit static rules for evasion.
Collectively, these Android XAI approaches underscore the interplay between interpretability and the fluid nature of the mobile ecosystem, where APIs and permissions often shift. Next, we examine how hardware-based XAI approaches address embedded systems and HPC-driven malware detection.